Prove that the sequence (fn(x)), where fn(x)= nx/(1+ nx^2) is not uniformly convergent in [-2,2]

1
Expert's answer
2022-06-07T15:20:09-0400

PROOF

Let "x=0," then "f_{n}(0)=0" for all "n\in\N" and "\lim_{n\rightarrow\infty}f_{n}(0)=0" . If "x\neq0" ,then "\lim_{ n \rightarrow\infty} f_{n}(x)=\lim_{n\rightarrow\infty} f_{n}(x)=\lim_{n\rightarrow\infty}\frac{nx} {1+nx^{2}}=\lim_{n\rightarrow \infty}\frac{nx} {n(\frac{1}{n}+ x^{2})}=\lim_{n\rightarrow\infty}\frac{ x} { (\frac{1}{n}+ x^{2})}=\\=\frac{x}{x^{2}}=\frac{1}{x} ."

Thus "f_{n}(x)" converges pointwise to

"f(x)= \begin{cases}
 0& \text{ if } x=0 \\ 
 \frac{1}{x}& \text{ if } x\in [-2,0)\bigcup (0,2]
\end{cases}" .

Note that the functions "f_{n}(x)" are continuous on the segment "[-2,2]" , the function "f" is discontinuous. Therefore, by Theorem 1 (see bellow) "f_{n}(x)" does not converge uniformly on "[-2,2]" .


THEOREM 1. Let "(f_{n}(x))" be a sequence of function on "A" converging uniformly to "f" on "A" .If each of "f_{n}" is continuous on "A" , then "f" is continuous on "A"